OpenStudy (anonymous):

lim sin(1/n)*n n->infinity

OpenStudy (ash2326):

put 1/n =y y-> 0 \[\lim_{y\to 0}\frac{\sin y}{y}\]

OpenStudy (anonymous):

After that

OpenStudy (anonymous):

Should I use L'Hospital rule @ash2326

OpenStudy (raden):

= 1

OpenStudy (anonymous):

Well thanx guys.

OpenStudy (ash2326):

it's a standard limit \[\lim_{x\to 0}\frac{\sin x}{x}=1\]

OpenStudy (anonymous):

oh ya.

OpenStudy (anonymous):

What if it was sin(x/n)*n where x is any integer.

OpenStudy (anonymous):

As n--->infinity

OpenStudy (ash2326):

Let's work with that \[\lim_{n\to \infty}{\sin x/n}\times n\] n=1/y y->0 so we have \[\lim_{y\to 0}\frac{\sin xy}{y}\] multiply numerator and denominator by x \[\lim_{y\to 0}x\frac{\sin xy}{xy}\] \[x\lim_{y\to 0}\frac{\sin xy}{xy}\] \[x\times 1\] \[x\]

OpenStudy (anonymous):

Sorry but I think it is x*pi/180

OpenStudy (anonymous):

Never mind I think I got it

OpenStudy (ash2326):

if x is in degrees, then we'd do that

OpenStudy (anonymous):

That depends on what x is.

OpenStudy (anonymous):

Thanx @ash2326